OpenStudy (anonymous):

Using the intermediate value theorem show that sin((pi x)/2)=1-x has a solution in (0,1) Help!!!!!!

OpenStudy (anonymous):

I tried plugging in 0 for x but I ended up with sin(0)=1 which is not true.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

start with \[f(x)=\sin(\frac{\pi}{2}x)+x-1\]

OpenStudy (misty1212):

plug in 0, plug in 1 one will be positive, one will be negative that means it has to be zero in that interval

OpenStudy (anonymous):

Thanks :))))))))))

OpenStudy (primeralph):

Yeah @aawowaa I think you need a better understanding of the theory.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]